OpenStudy (anonymous):

the function f is defined by f(x)=2x^2-1 find f(2y) .......can someone help me please will medal

OpenStudy (faiqraees):

you haven't written the correct question

OpenStudy (faiqraees):

Is it "find f(2y)"

OpenStudy (anonymous):

yes i changed it :)

OpenStudy (anonymous):

@FaiqRaees

OpenStudy (faiqraees):

Okay so in place of x write 2y and rewrite the function

OpenStudy (faiqraees):

\[\large\rm f(x)=2x^2-1 \]\[\large\rm f(x)=2(2y)^2-1 \]

OpenStudy (anonymous):

oh ok i wrote it wrong

OpenStudy (faiqraees):

No you didnt write it wrong. The equations I wrote are part of the solution

OpenStudy (anonymous):

ok so would you take 2y times 2y then times it by 2 or am I way off base

OpenStudy (anonymous):

@FaiqRaees

OpenStudy (faiqraees):

correct

OpenStudy (faiqraees):

\[\large\rm f(x)=(2*2y*2y)-1 \]

OpenStudy (anonymous):

so would it be f(x) =8y-1 @FaiqRaees

OpenStudy (faiqraees):

No y times y is y squared so 2y times 2y is 4 y squared and 4 y squared times 2 is 8 y squared

OpenStudy (anonymous):

oops i forgot my squared

OpenStudy (anonymous):

would I write it as f(2y)=8y^2-1

OpenStudy (anonymous):

@FaiqRaees

OpenStudy (faiqraees):

Perfect

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

@FaiqRaees

OpenStudy (faiqraees):

You are welcome